OpenStudy (anonymous):

One of my MyMathLab problems is asking me to solve a logarithmic problem. I'm working through the problem correctly, but my calculator isn't giving me the right answer when I plug in the log with the coefficient (I'm trying to find 1.71^3x=3, which I've been plugging into the calculator as log(3)/3log(1.71). Am I entering the coefficient of the log correctly? One of my MyMathLab problems is asking me to solve a logarithmic problem. I'm working through the problem correctly, but my calculator isn't giving me the right answer when I plug in the log with the coefficient (I'm trying to find 1.71^3x=3, which I've been plugging into the calculator as log(3)/3log(1.71). Am I entering the coefficient of the log correctly? @Mathematics

OpenStudy (zarkon):

this\[1.71^{3x}=3\]?

OpenStudy (anonymous):

Exactly. I guess my question is, how do you enter a coefficient of a logarithm into a TI-83 plus calculator? Like I said, when I plug it in as "3log(1.71)" it's not giving me the correct answer

OpenStudy (zarkon):

\[\frac{\ln(3)}{3\ln(1.71)}\]

OpenStudy (zarkon):

\[\approx.68258829717\]

OpenStudy (anonymous):

@Zarkon: can you pls help me in my question!!!!

OpenStudy (anonymous):

Awesome, thank you! I need to fix something with my calculator settings, because when I plugged that in it gave me .1964660699...